Dr. Bryan Carpenter

Department of Electronics and Computer Science University of Southampton Highfield Southampton S09 5NH

Tel: 0703 593081

email: dbc@uk.ac.soton.ecs

August 24, 1994

Dr. Pieter Hartel, Department of Computer Systems, University of Amsterdam, Kruislaan 403, 1098 SJ Amsterdam, The Netherlands

Dear Pieter,

I’m enclosing my comments on the paper “Comparing Id and Haskell in a Monte Carlo Photon Transport Code” by Jeffrey Hammes, Olaf Lubeck, Wim Boehm.

I think the paper is reasonably well written, and that it’s an interesting thing to try, and a modified version of the paper ought to be acceptable.

(I have a feeling that a lot of their work on the Id code isn’t about functional programming, as such, but I’m not sure what can be done about that).

Yours sincerely,